OpenStudy (anonymous):

simplify please (sec(x))^(2)−1 simplify please (sec(x))^(2)−1 @Mathematics

OpenStudy (anonymous):

(tan(x))^2

OpenStudy (anonymous):

oh wo zekarias thanks for the help

OpenStudy (anonymous):

wow*

OpenStudy (anonymous):

might you be able to help me out with another problem

OpenStudy (anonymous):

off course

OpenStudy (amistre64):

these are just the pythag identities \[sin^2+cos^2=1\] \[\frac{sin^2}{cos^2}+\frac{cos^2}{cos^2}=\frac{1}{cos^2}\] \[tan^2+1=sec^2\]

OpenStudy (anonymous):

If the point (2, -2) is on the terminal side of the angle θ, find each of sin(θ),cos(θ),tan(θ)

OpenStudy (anonymous):

oh ok thanks amistre...hey i apologize about what happened last time

OpenStudy (amistre64):

'sok, it happens :)

OpenStudy (anonymous):

zekarias are you there

OpenStudy (anonymous):

\[\sin \theta=-1/\sqrt{2}\]\[\cos \theta=1/\sqrt{2}\]\[\tan \theta=-1\]

OpenStudy (anonymous):

wow thanks how did you figure that...i've been trying to do it for hours. also how can I become a fan of you...you have really helped me